OpenStudy (anonymous):

Pre calculus math question? help please? The value 3pi/4 is a solution for the equation 3square root of 2 cos theta +2=-1 A. True B. False

OpenStudy (anonymous):

\[\bf 3\sqrt{2}\cos(\theta)+2=-1\]Is this what you mean? @StudentLearner

OpenStudy (anonymous):

yes @genius12

OpenStudy (anonymous):

Rearrange the equation to get:\[\bf \cos(\theta)=\frac{ -3 }{ 3\sqrt{2} }=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

so that means it would be false? :)

OpenStudy (jdoe0001):

look at your Unit Circle

OpenStudy (jdoe0001):

\(\frac{3\pi}{4}\) would be in the 2nd Quadrant

OpenStudy (anonymous):

We know from our common trig values that:\[\bf \cos \left( \frac{\pi}{4} \right)=\frac{\sqrt{2}}{2}\]But that angle is in the first quadrant. Cosine is negative in the second and third quadrants. Hence the reference angles that will result in a negative value instead of a positive value will be:\[\bf x= \frac{3\pi}{4} \ or \ x=\frac{5\pi}{4}\] @StudentLearner

OpenStudy (anonymous):

so it is true :D thanks guys